Your company is considering the purchase of a new machine. After 4 years of operation, you would sell the machine for a salvage value of $71297. However, at that time the machine would not have been depreciated to a value of 0, but have a remnant book value of $23607. The operation of the machine requires additional NOWC of $44180, which would not change throughout the project. The firm's corporate tax rate is 40%. What is the project's Terminal Cash Flow?

Enter your answer in dollars, rounded to 2 decimals, without the dollar sing. So, if your answer is 12,345.6789, just enter 12345.68.
Answer #1

TERMINAL CASH FLOW = AFTER TAX SALVAGE VALUE OF MACHINE + RECOVERY OF WORKING CAPITAL

AFTER TAX SALVAGE VALUE OF MACHINE = SALVAGE VALUE - TAX ON PROFIT ON SALE OF MACHINE

AFTER TAX SALVAGE VALUE OF MACHINE = 71297 - 40%*(71297 - 23607) = 52221

RECOVERY OF WORKING CAPITAL = 44180

TERMINAL CASH FLOW = 52221 + 44180 = 96401

ANSWER : 96401.00 (Thumbs up please)